Encyclopedia Astronautica
2007.12.14 - ISS On-Orbit Status 12/14/07


The CDR and FE-2 conducted a one-hour review of an uplinked procedures briefing package for the US EVA-13 next week (12/18), covering topics like egress plan, timeline ordering of tasks, translation/fairleads/tether plan, hazards, and ingress plan.

After wakeup and before breakfast, FE-2 Dan Tani again accessed the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and completing questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for later downlink. (To monitor the crewmember's sleep/wake patterns and light exposure, Dan wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

Also upon wakeup, FE-1 Yuri Malenchenko terminated his fourth MBI-12 SONOKARD experiment session by taking the recording device from his SONOKARD sports shirt pocket and later copying the measurements to the RSE-MED laptop for subsequent downlink to the ground. (SONOKARD objectives are stated to (1) study the feasibility of obtaining the maximum of data through computer processing of records obtained overnight, (2) systematically record the crewmember's physiological functions during sleep, (3) study the feasibility of obtaining real-time crew health data. Investigators believe that contactless acquisition of cardiorespiratory data over the night period could serve as a basis for developing efficient criteria for evaluating and predicting adaptive capability of human body in long-duration space flight.)

The FE-2 again activated the VDS MPC (Video Distribution System/Multi-Purpose Converter) with its four downlinks to allow the ground to conduct HDTV (high-definition TV) playback and downlink operations. Later today (~1:40pm EST), the MPC was powered off again.

Tani then assisted the ground in ongoing troubleshooting of the transmission & downlinking of analog (as opposed to digital) video signals from the Russian Segment (RS) via the MPEG-2 (Moving Pictures Expert Group 2) encoder and Ku-band in 'streaming video' packets over the U.S. OpsLAN. (Prior end-to-end video test results had not been as expected, and the evaluation of downlinked log files continues. For today's troubleshooting, Tani set up the software application for PAL (Russian)-to-NTSC (US) format conversion on an A31p laptop, connected cables and started the laptop. After the test, the A31p was left on for eventual downlinking of more log files.)

Malenchenko completed the periodic switchover of the Russian STTS telephone/telegraph subsystem to an alternate string, today to the primary string after its operation on the backup string. (The "Voskhod-M" STTS enables telephone communications between the SM, FGB, DC1 and U.S. segment (USOS), and also with users on the ground over VHF channels selected by an operator at an SM comm panel, via STTS antennas on the SM's outside. There are six comm panels in the SM with pushbuttons for accessing any of three audio channels, plus an intercom channel. Other modes of the STTS include telegraphy (teletype), EVA voice, emergency alarms, Packet/Email, and TORU docking support.)

Dan Tani had about 4 hrs scheduled for auditing and consolidating photo/TV items in their CTBs (Cargo Transfer Bags), including a large number of size AA batteries. (This task has up to now been on the U.S. 'job jar' task list for Dan and Peggy and was now hard scheduled due to crewtime availability.)

The CDR and FE-2 conducted a one-hour review of an uplinked procedures briefing package for the US EVA-13 next week (12/18), covering topics like egress plan, timeline ordering of tasks, translation/fairleads/tether plan, hazards, and ingress plan. Later (~7:00am EST) Peggy and Dan tagged up with ground specialists to discuss particulars. (The spacewalk of ~6.5 hrs duration has two major objectives: (1) Inspection of 1A BGA (Beta Gimbal Assembly) and BMRMM (Bearing Motor Roll Ring Module, 'broom'), including assisting fault search by the ground by disconnecting/reconnecting cables and possibly performing an R&R (removal & replacement) of the 1A ECU (Electronic Control Unit) on the S4 truss; (2) inspection and photo documentation of the Stbd SARJ (Solar Alpha Rotary Joint), including temporary removal of protective MLI covers (8 double-wide, 12 single-wide, plus 2 DLA/Drive Lock Assembly covers), debris removal, DLA inspection, finally unbolting (3 bolts) & removing TBA-5 (Trundle Bearing Assembly #5) for return to Earth.)

The FE-1 meanwhile worked in the RS on the line 3 BRPK-2 Air/Liquid Condensate Separator of the SRV-K2M Condensate Water Recovery System, removing the mechanical filter cartridge from the separator's transfer unit.

Whitson took air samples for the periodic (weekly) atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the hand-held CSA-O2 (Compound Specific Analyzer-Oxygen sensor) and CDMK (CO2 Monitoring Kit). Batteries were to be replaced if necessary. (Purpose of the 15-min activity is to trend with MCA (Major Constituents Analyzer), i.e., to correlate the hand-held readings with MCA measurements. O2 sensors checks used #1042, #1063, #1052, #1041. CO2 sensors used are #1013, #1058, #1045. CO2 level in Lab and SM was ~0.44%.)

With the BCAT-3 (Binary Colloidal Alloy Test-3) science payload running by itself in Node-2 since yesterday, FE-2 Tani began a daily status check on the A31p laptop controlling the EarthKAM digital still camera, verifying on the last image taken that image focus and alignment remain in check. (The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour for the next few days.)

Yuri handled the daily IMS maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Peggy and Yuri each had another ~4.5 hrs reserved for stowing U.S. and Russian trash cargo on the Progress 26P cargo ship-turned-trash can, currently docked at the DC-1 nadir port. (Discarded hardware, as specified on uplinked lists, includes such items as Russian food containers (U.S. food containers being recycled on the ground), food waste bags, Penguin-3 suits, TVIS malfunction kit & harnesses, payload containers, one old 760XD laptop, socks, coveralls, wipes & wet towels, medical kits, used & expired alkaline batteries, 35-mm film cassettes, hoses & adapters, cartridge belts with tools, foam pieces, etc.)

Dan conducted the periodic (every two weeks) 10-min inspection of the RED (Resistive Exercise Device) canister cords and accessory straps as well as the canister bolts for re-tightening if required.

The crewmembers performed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Peggy copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

A new task added to the voluntary 'job jar' list for the CDR and FE-2 calls for an audit of rubber gloves on board, to report the quantity of available clean room rubber gloves and Nitrile gloves stowed in a CTB and in the ASU toilet facility.

At ~2:30am EST, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU), including Shift Flight Director (SRP), at TsUP via S-band/audio, phone-patched from Houston and Moscow.

At ~4:00am, Yuri linked up with TsUP stowage specialists via S-band to conduct the weekly IMS tagup, discussing stowage issues and equipment locations.

At ~8:20am, the FE-1 also powered up the SM (Service Module)'s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, and power supply), to conduct, at 8:25am, a ham radio exchange with students and faculty at Kursk State Polytechnic University who actively participated in amateur sessions from Kursk, radio expeditions from Sochi (Dagomys) and from the peak of Mt. Elbrus with Expedition 15 Commander Fyodor Yurchikhin.

At ~2:00pm, the crew conducted their sixth weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio. (S/G-2 (Space-to-Ground 2) phone patch via SSC-10 (Station Support Computer 10)).

KURS Antenna Test: Yesterday, a ground-commanded test of the KURS system was performed over RGS (Russian Ground Sites) from the DC1 Docking Compartment side in preparation for Progress M-62/27P docking on 12/26. Test results for string 1 of the KURS hardware were nominal, but results for string 2 were off-nominal. Specialists at TsUP-Moscow believe the off-nominal results may be due to a loss of VHF (Very High Frequency)-1 communication and the telemetry stream from RGS 27 (BRL). TsUP will decide if a retest on string 2 is required.

TORU Test: In preparation for 27P arrival, Moscow has scheduled a test of the TORU manual teleoperator control system in the SM for 12/17.

26P Rodnik Tank: Since 9/4, the expulsion bladder in Progress 26P's Rodnik tank 1 has been suspected of a small leak, possibly rendering the tank unfit for the upcoming urine transfer from the SM. TsUP this week learned from the supplier that the bladder is acceptable for urine transfer provided the KN1 valve is closed during the pump-over, currently scheduled for 12/17 (Monday).

MT Translation: At 11:41am-1:41pm the MT (Mobile Transporter) was remotely driven from WS7 (Worksite 7) to WS4, in order to protect the MT TUS (Trailing Umbilical System) cable from MMOD (micrometeoroid/orbital debris) damage. (In pre-configuring the RS systems for the thruster-disabled period duringt the translation (to reduce loads), a software program uplinked was inadvertently executed, configuring the Russian SUDN (Motion Control & Navigation System) such that both it and the U.S. GNC (Guidance, Navigation & Control) systems were briefly in control of ISS attitude. Action to reconfigure into a structurally safe mode was taken immediately. The rail cart will be returned to WS4 shortly before 1E arrival. Before the translation, the rail path was surveyed with the SSRMS (Space Station Remote Manipulator System) vidcams for obstructions.)

CEO (Crew Earth Observation) photo targets uplinked for today were Somalia Coast, Somalia (weather was predicted to be clear over Somalia for photography of coastal dune and vegetation patterns. Collection of baseline imagery during the current La Nina conditions will be useful for comparison with the next El Nino event. Looking to the right of track as ISS proceeded NE parallel to the coastline; overlapping, along-track frames were recommended), and Gulf of Fonseca, Honduras (scattered clouds were predicted over the Gulf of Fonseca, but there should have been opportunities for mapping photography of the site. Overlapping frames of the coastline were requested. Repeat imagery of the site will allow for assessment of land use and land cover change on the coastal environment over time).

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use